# Count on Your Elders: Laplace vs Gaussian Noise

### v2

@@ 4.1 Privacy
We proceed by proving the privacy of Algorithm 3, the privacy of which implies the privacy of Algorithm 2 via Proposition 4.3. The privacy proof is standard and follows from identifying that if all $\widehat{\Sigma}(I)$ on line 4 are released privately, then the remaining computations are post-processing of that release. For neighbouring inputs $\mathbf{x}^{\prime}\sim\mathbf{x}$, let $i$ be the index at which they differ. Let $\mathbf{y}\in\mathbb{R}^{|\mathcal{T}_{k,h}|}$ be a vector where each entry is identified by a unique interval from $I\in\mathcal{T}_{k,h}$ equal to $\Sigma(I)$ given $\mathbf{x}$. Define $\mathbf{y}^{\prime}$ analogously for $\mathbf{x}^{\prime}$. Since any $i\in[1,T]$ is included in exactly $h$ intervals $I\in\mathcal{T}$, intentionally excluding the root vertex which is not used, it follows that the sensitivity for releasing $\mathbf{y}$ is $\Delta_{1}=\left\lVert\mathbf{y}-\mathbf{y}^{\prime}\right\rVert_{1}=h$. As $\mathbf{y}$ is released on line 4 using the Laplace mechanism with appropriate scale, and all computations afterward are post-processing of $y$, Algorithm 3 satisfies $\varepsilon$-DP.
@@ 4.3 Space and time complexity
Each noise term used $\mathbf{z}_{p}$ used by Algorithm 2 will only be used for outputs in a time interval $[t_{1},t_{2}]\subseteq[1,T]$.
@@ 5 Laplace vs Gaussian noise
Furthermore, we adapt the Laplace mechanism to work with $\ell_{2}$-sensitivity and show that it provides a meaningful $(\varepsilon,\delta)$-DP guarantee with variance close to that of the Gaussian mechanism. For proofs not included we refer the reader to Appendix D.
@@ 5.2 Laplace noise scaled to $\ell_{2}$-sensitivity
We next consider how we may extend the Laplace mechanism with an approximate DP guarantee similar to the Gaussian mechanism, i.e., with noise scaled to the $\ell_{2}$-sensitivity. Since the proof of this is straightforward using known results we suspect this fact may be known, but we are not aware of it appearing in the literature.
@@ 6 Discussion
While there are no known fundamental reasons ruling out asymptotic improvements in the mean squared error for $\varepsilon$-DP, it might be possible to match the $\Omega(\log(T)^{2})$ lower bound of [24], there have been no asymptotic improvements since the binary tree mechanism [16, 5]. This state of affairs motivates exploring how much leading constants can be improved using the factorization mechanism framework. Algorithm 2 falls into the framework, is easy to implement, efficient, and offers the best known error-scaling to date.
Our paper also reveals that using Laplace noise can improve known “optimal” factorization mechanisms using Gaussian noise under rather weak assumptions on $\delta$. An interesting future direction is to investigate, for fixed $\varepsilon,\delta$ and a given factorization: which noise distribution achieves optimal error? It seems unlikely that Laplace or Gaussian noise is always best – indeed, Vinterbo [37] has results of this flavor in more restricted settings.

### v3

@@ 4.1 Privacy
We proceed by proving the privacy of Algorithm 3, the privacy of which implies the privacy of Algorithm 2 via Proposition 4.3. The privacy proof is standard and follows from identifying that if all $\widehat{\Sigma}(I)$ on line 4 are released privately, then the remaining computations are post-processing of that release. For neighboring inputs $\mathbf{x}^{\prime}\sim\mathbf{x}$, let $i$ be the index at which they differ. Let $\mathbf{y}\in\mathbb{R}^{|\mathcal{T}_{k,h}|}$ be a vector where each entry is identified by a unique interval from $I\in\mathcal{T}_{k,h}$ equal to $\Sigma(I)$ given $\mathbf{x}$. Define $\mathbf{y}^{\prime}$ analogously for $\mathbf{x}^{\prime}$. Since any $i\in[1,T]$ is included in exactly $h$ intervals $I\in\mathcal{T}$, intentionally excluding the root vertex which is not used, it follows that the sensitivity for releasing $\mathbf{y}$ is $\Delta_{1}=\left\lVert\mathbf{y}-\mathbf{y}^{\prime}\right\rVert_{1}=h$. As $\mathbf{y}$ is released on line 4 using the Laplace mechanism with appropriate scale, and all computations afterward are post-processing of $y$, Algorithm 3 satisfies $\varepsilon$-DP.
@@ 4.3 Space and time complexity
Each noise term $\mathbf{z}_{p}$ used by Algorithm 2 will only be used for outputs in a time interval $[t_{1},t_{2}]\subseteq[1,T]$.
@@ 5 Laplace vs Gaussian noise
Furthermore, we adapt the Laplace mechanism to work with $\ell_{2}$-sensitivity and show that it provides a meaningful $(\varepsilon,\delta)$-DP guarantee with variance close to that of the Gaussian mechanism.
@@ 5.2 Laplace noise scaled to $\ell_{2}$-sensitivity
We next consider how we may extend the Laplace mechanism with an approximate DP guarantee similar to the Gaussian mechanism, i.e., with noise scaled to the $\ell_{2}$-sensitivity. Since the proof of this is straightforward using known results we suspect this fact may be known, but we are not aware of it appearing in the literature. The technique for extending the Laplace mechanism to approximate DP relies on the following heterogeneous composition theorem [30].
Lemma 5.4 (Simplified version of Theorem 3.5 in [30]).
Using Lemma 5.4, we can derive the following approximate DP guarantee for the $\ell_{2}$ Laplace mechanism.
@@
Proof 5.6.
For $x_{1},\dots,x_{j}\in(0,1)$, since $\frac{e^{x_{i}}-1}{e^{x_{i}}+1}\leq x_{i}/2$ it follows from Lemma 5.4 that $(\tilde{\varepsilon},\delta)$-differential privacy is also satisfied for
@@
Proof 5.8.
Finally we prove Theorem 1.3 from the introduction.
Proof 5.9 (Proof of Theorem 1.3).
@@
6 Lower bound for continual observation on random inputs
In this section we consider continual observation in the “offline” setting where a mechanism has access to the entire input and has to output a differentially private approximation of all prefix sums. This of course implies a lower bound for the “online” setting. We show Theorem 1.4 which achieves the same lower bound that was previously known for worst-case inputs, in the setting where the input is a random string.
Proof 6.1 (Proof of Theorem 1.4).
We will prove the theorem via a packing argument applied to a carefully constructed collection of random strings. Without loss of generality, let $B=\sqrt{T}$ be an integer divisible by $4$, $k\leq T^{1/5}$ be an even integer, and $T$ be large enough such that $k\leq B/4$. Also divide the timeline $[1,T]$ into blocks of length $B$: $B_{1}=[1,B],B_{2}=[B+1,2B],\dots,B_{m}=[T-B+1,T]$, $m=T/B$. For an arbitrary string $\mathbf{y}\in\{0,1\}^{T}$, we use the notation $\mathbf{y}_{B_{i}}\in\{0,1\}^{B}$ for the substring of $\mathbf{y}$ formed from the indices in $B_{i}$. We use $\mathrm{err}(\cdot)$, taking as input the output of a mechanism, to denote the (random) maximum additive error over all the outputs. The argument is outlined below:
  • Construct random strings $\mathbf{x}^{(0)},\dots,\mathbf{x}^{(m)}$ that are near-neighboring and have small total variation distance to random strings from $\mathcal{U}$.
  • Argue that $\mathcal{M}$ assumes additive error $\leq\alpha$ on these inputs with constant probability by virtue of the small total variation distance.
[…]
The intuition behind the argument is that an accurate mechanism $\mathcal{M}$ on $\mathcal{U}$ will also be accurate on our random (but highly correlated) strings $\mathbf{x}^{(0)},\dots\mathbf{x}^{(m)}$. Being accurate on these strings implies the ability to distinguish between them. However, since $\mathcal{M}$ satisfies $\varepsilon$-DP, distinguishing between the inputs with good confidence is impossible, implying a lower bound on the additive error $\alpha$.
To this end, we define the random variables $X^{(i)}_{j}=\left\lVert\mathbf{x}^{(i)}_{B_{j}}\right\rVert_{1}=\sum_{\ell\in B_{j}}\mathbf{x}^{(i)}_{\ell}$, i.e., as the sum of 1s in the block $B_{j}$ of $\mathbf{x}^{(i)}$. With the argument outlined, consider the random strings $\mathbf{x}^{(0)},\mathbf{x}^{(1)},\dots,\mathbf{x}^{(m)}$ formed by the following process:
  1. 1.
    Draw $\mathbf{x}^{(0)}$ from $\mathcal{U}$ conditioned on $X^{(0)}_{j}\in[\frac{B}{4},\frac{3B}{4}]$ for all $j\in[m]$. Call this distribution $\mathcal{D}^{(0)}$.
  2. 2.
    For $i\in[m]$ derive $\mathbf{x}^{(i)}$ from $\mathbf{x}^{(0)}$ by choosing $k$ 0s in $B_{i}$ uniformly at random and flipping them to 1s. Call these distributions $\mathcal{D}^{(i)}$.
We next define the algorithm $\mathrm{Alg}(\mathbf{y},\mathbf{y}^{\prime})$ outputting the difference between private sums on two arbitrary inputs $\mathbf{y},\mathbf{y}^{\prime}\in\{0,1\}^{T}$. Formally, on receiving $\mathbf{y}$ and $\mathbf{y}^{\prime}$, it runs $\mathcal{M}$ on each, producing outputs $\mathbf{a}_{1},\dots,\mathbf{a}_{T}$ and $\mathbf{a}_{1}^{\prime},\dots,\mathbf{a}_{T}^{\prime}$ respectively, and at time $t$, the algorithm outputs $\mathrm{Alg}(\mathbf{y},\mathbf{y}^{\prime})_{t}=\mathbf{a}_{t}-\mathbf{a}_{t}^{\prime}$. We will show, step-by-step, that the accuracy guarantee of $\mathcal{M}$ on $\mathcal{U}$ will imply distinguishing between our random strings $\mathbf{x}^{(0)},\dots,\mathbf{x}^{(i)}$.
Consider $\mathrm{Alg}(\mathbf{x}^{(i)},\mathbf{x}^{(0)})$ for $i\in[1,m]$. Note that the true prefix sums of $\mathbf{x}^{(i)}$ and $\mathbf{x}^{(0)}$ at the end of each block will be identical up until block $B_{i}$, at the end of which they will differ by exactly $k$. We leverage this observation by introducing the following disjoint events $E_{j}$. For $j\in[1,m]$ and arbitrary input strings $\mathbf{y}$, $\mathbf{y}^{\prime}$, we define $E_{j}$ as the event that the following two statements are true: (1) for all $\ell\in[1,j-1]:\mathrm{Alg}(\mathbf{y},\mathbf{y}^{\prime})_{\ell B}\leq k/2$, and (2) $\mathrm{Alg}(\mathbf{y},\mathbf{y}^{\prime})_{jB}>k/2$. In other words, $E_{j}$ is the event that, when tracking only the outputs at the end of blocks, $\mathrm{Alg}(\mathbf{y},\mathbf{y}^{\prime})$ outputs a value greater than $k/2$ for the first time at the end of block $B_{j}$. Set $\alpha=k/4$. By our construction, if $\mathcal{M}(\mathbf{x}^{(0)})$ and $\mathcal{M}(\mathbf{x}^{(i)})$ each assumes additive error at most $\alpha$ when running $\mathrm{Alg}(\mathbf{x}^{(i)},\mathbf{x}^{(0)})$, then $\mathrm{Alg}(\mathbf{x}^{(i)},\mathbf{x}^{(0)})$ has additive error at most $22\alpha=k/$ 2, and $E_{i}$ happens with probability $1$. We thus have
$\Pr[\mathrm{Alg}(\mathbf{x}^{(i)},\mathbf{x}^{(0)})\in E_{i}]\geq\Pr[\mathcal{M}(\mathbf{x}^{(i)})\leq\alpha]\cdot\Pr[\mathcal{M}(\mathbf{x}^{(0)})\leq\alpha]\enspace.$
We want to argue that $\mathcal{M}$ is accurate on $\mathbf{x}^{(0)},\dots,\mathbf{x}^{(m)}$, with constant probability for each input taken separately, and we will argue via the total variation distance between each of these random strings, and the random string $\mathbf{x}\sim\mathcal{U}$. In general, for a (possibly randomized) algorithm $\mathcal{A}:S\to\mathbb{R}$, distributions $\mathcal{D},\mathcal{D}^{\prime}$ over $S$ with total variation distance $\operatorname{TV}(\mathcal{D},\mathcal{D}^{\prime})\leq\Delta$, and an event $E$ over the output, we have that if $\Pr_{\mathcal{A},x\sim\mathcal{D}}[\mathcal{A}(x)\in E]\geq p\geq 0$, then $\Pr_{\mathcal{A},x\sim\mathcal{D}^{\prime}}[\mathcal{A}(x)\in E]\geq p-\Delta$. This follows from the observation that, in the extreme case, $\Delta$ probability mass can be shifted from inputs $x\in S$ where $\Pr_{\mathcal{A}}[\mathcal{A}(x)\in E]=1$ to inputs $x^{\prime}\in S$ where $\Pr_{\mathcal{A}}[\mathcal{A}(x^{\prime})\in E]=0$. If we manage to show that, for $i\in[0,m]$, $\operatorname{TV}(\mathbf{x},\mathbf{x}^{(i)})=o(1)$, then we can lower bound $\Pr[\mathrm{Alg}(\mathbf{x}^{(i)},\mathbf{x}^{(0)})\in E_{i}]$ by a positive constant. We derive bounds on these total variation distances in the next two lemmas.
Lemma 6.2.
Let $\mathbf{x}\sim\mathcal{U}$ and $\mathbf{x}^{(0)}\sim\mathcal{D}^{(0)}$. Then $\operatorname{TV}(\mathbf{x},\mathbf{x}^{(0)})\leq\Delta$ where $\Delta=2\sqrt{T}\exp(-\sqrt{T}/8)$
Proof 6.3.
Let $S$ be the support of $\mathcal{U}$ and $S^{(0)}\subseteq S$ the support of $\mathcal{D}^{(0)}$. We have that $X\sim\mathrm{Bin}(B,1/2)$ describes the number of 1s in an arbitrary block $B_{i}$ of $\mathbf{x}$. Hoeffding’s inequality gives $\Pr[\frac{B}{4}\leq X\leq\frac{3B}{4}]\geq 1-2\exp(-B/8)$ whereby it follows that $\Pr[\mathbf{x}\in S^{(0)}]\geq\big{(}1-2\exp(-B/8)\big{)}^{T/B}$. Observing that $\mathbf{x}^{(0)}$ is uniformly distributed over $S^{(0)}\subseteq S$, we proceed to directly compute the total variation distance:
$\operatorname{TV}(\mathbf{x},\mathbf{x}^{(0)})$ $=\frac{1}{2}\sum_{\mathbf{s}\in S}|\Pr[\mathbf{x}=\mathbf{s}]-\Pr[\mathbf{x}^{(0)}=\mathbf{s}]|$
$=\frac{1}{2}|S^{(0)}|\bigg{(}\frac{1}{|S^{(0)}|}-\frac{1}{|S|}\bigg{)}+\frac{1}{2}(|S|-|S^{(0)}|)\cdot\frac{1}{|S|}=1-\frac{|S^{(0)}|}{|S|}=1-\Pr[\mathbf{x}\in S^{(0)}]$
$\leq 1-\big{(}1-2\exp(-B/8)\big{)}^{T/B}\leq\frac{2T\exp(-B/8)}{B}=2\sqrt{T}\exp(-\sqrt{T}/8)\enspace,$
where the last inequality is using Bernoulli’s inequality and holds for $T$ larger than an absolute constant.
We prove that the remaining random strings $\mathbf{x}^{(1)},\dots,\mathbf{x}^{(m)}$ are closely distributed to strings in $\mathcal{U}$ next.
Lemma 6.4.
Let $\mathbf{x}\sim\mathcal{U}$ and $\mathbf{x}^{(i)}\sim\mathcal{D}^{(i)}$ for $i\in[m]$. Then $\operatorname{TV}(\mathbf{x},\mathbf{x}^{(i)})\leq\Delta^{\prime}$ where $\Delta^{\prime}=10\cdot T^{-1/20}$.
Proof 6.5.
As the total variation distance satisfies the triangle inequality, we can write
$\operatorname{TV}(\mathbf{x},\mathbf{x}^{(i)})\leq\operatorname{TV}(\mathbf{x},\mathbf{x}^{(0)})+\operatorname{TV}(\mathbf{x}^{(0)},\mathbf{x}^{(i)})\enspace.$
The first term we have already computed in Lemma 6.2. To bound the second term, we will use a property of total variation distance: it cannot increase if one applies the same transformation to both arguments. Let $f:[m]\times[0,B]\to\mathcal{F}$ be a mapping to a family of distributions where $f(i,\ell)$ denotes the distribution described by a process where (1) a string $\mathbf{y}\sim\mathcal{D}^{(0)}$ is sampled and then (2) $\mathbf{y}_{B_{i}}$ is overwritten by a uniformly sampled string in $\{0,1\}^{B}$, conditioned on it having $\ell$ 1s. For $i\in[m]$, we claim that ($f(i,X^{(i)}_{i})$, $\mathcal{D}^{(i)})$ and $(f(i,X^{(0)}_{i}),\mathcal{D}^{(0)})$ are pairs of identical distributions. To see this for the first pair of distributions, consider $\mathbf{y}\sim f(i,X^{(i)}_{i})$. By definition $\mathbf{y}_{B_{j}}$ is identically distributed to $\mathbf{x}^{(i)}_{B_{j}}$ for $j\neq i$, and so the only special case to consider is $j=i$. For this case, we note that $X^{(i)}_{i}$ is defined as the number of 1s in $\mathbf{x}^{(i)}_{B_{i}}$, and that symmetry implies that any substring with $\ell$ 1s is equally likely, thus $\mathbf{y}_{B_{i}}$ must be identically distributed to $\mathbf{x}^{(i)}_{B_{i}}$ as well, and we are done. An analogous argument can be used for showing that $f(i,X^{(0)}_{i})$ and $\mathcal{D}^{(0)}$ are identically distributed.
We thus have $\operatorname{TV}(\mathbf{x}^{(0)},\mathbf{x}^{(i)})=\operatorname{TV}(f(i,X^{(0)}_{i}),f(i,X^{(i)}_{i}))\leq\operatorname{TV}(X^{(0)}_{i},X^{(i)}_{i})$, and so our problem reduces to bounding this quantity.
$\operatorname{TV}(X^{(0)}_{i},X^{(i)}_{i})$ $=\frac{1}{2}\sum_{\ell=0}^{B}\lvert\Pr[X^{(0)}_{i}=\ell]-\Pr[X^{(i)}_{i}=\ell]\rvert$
$=\frac{1}{2}\sum_{\ell=0}^{B}\lvert\Pr[X^{(0)}_{i}=\ell]-\Pr[X^{(0)}_{i}=\ell-k]\rvert$
$=\frac{1}{2}\bigg{(}\sum_{\ell=B/4}^{B/4+k-1}\Pr[X^{(0)}_{i}=\ell]+\sum_{l=3B/4+1-k}^{3B/4}\Pr[X^{(0)}_{i}=\ell]$
$+\sum_{\ell=B/4+k}^{3B/4}\lvert\Pr[X^{(0)}_{i}=\ell]-\Pr[X^{(i)}_{i}=\ell]\rvert\bigg{)}$
$\leq k\cdot\Pr[X^{(0)}_{i}=B/4+k]+\frac{1}{2}\sum_{\ell=B/4+k}^{3B/4}\lvert\Pr[X^{(0)}_{i}=\ell]-\Pr[X^{(0)}_{i}=\ell-k]\rvert$
The last step follows from the symmetry of $X^{(0)}_{i}$ : $\Pr[X^{(0)}_{i}=B/2+\ell]=\Pr[X^{(0)}_{i}=B/2-\ell]$, and that the probability is greater for values closer to $B/2$. Observe that the first $B/4-k/2$ terms in the second sum of the last step are equal to the last $B/4-k/2$ terms by the symmetry of the distribution around $B/2$ (for $\ell=B/2+k/2$, the term in the sum is equal to zero). Continuing the derivation and leveraging this symmetry, we get
$\operatorname{TV}(X^{(0)}_{i},X^{(i)}_{i})$ $\leq k\cdot\Pr[X^{(0)}_{i}=B/4+k]+\sum_{\ell=B/4+k}^{B/2+k/2-1}(\Pr[X^{(0)}_{i}=\ell]-\Pr[X^{(0)}_{i}=\ell-k])$
$=k\cdot\Pr[X^{(0)}_{i}=B/4+k]$
$+\sum_{\ell=0}^{k-1}(\Pr[X^{(0)}_{i}=B/2-k/2+\ell]-\Pr[X^{(0)}_{i}=B/4+\ell])$
$\leq 2k\cdot\Pr[X^{(0)}_{i}=B/2]$
where the first equality follows from the sum telescoping and the second inequality follows from bounding each term in the sum. Continuing, letting $X\sim\mathrm{Bin}(B,1/2)$, we have that
$2k\cdot\Pr[X^{(0)}_{i}=B/2]$ $=2k\cdot\Pr\bigg{[}X=B/2\,\bigg{|}\,\frac{B}{4}\leq X\leq\frac{3B}{4}\bigg{]}=\frac{2k\cdot\Pr[X=B/2]}{1-2\Pr[X<B/4]}$
$\leq\frac{2k\cdot\Pr[X=B/2]}{1-2\exp(-B/8)}\leq 6k\cdot\Pr[X=B/2]=6k\cdot 2^{-B}\binom{B}{B/2}$
$\leq 6k\cdot 2^{-B}\cdot\frac{\sqrt{2}\cdot 2^{B}}{\sqrt{\pi B}}=\frac{6\sqrt{2}}{\sqrt{\pi}}\cdot\frac{k}{\sqrt{B}}=\frac{6\sqrt{2}}{\sqrt{\pi}}\cdot T^{-1/20}\enspace,$
where all inequalities are valid for $T$ greater than some absolute constant, and the last inequality is a variant of Stirling’s inequality. We finally get
$\operatorname{TV}(\mathbf{x},\mathbf{x}^{(i)})\leq\operatorname{TV}(\mathbf{x},\mathbf{x}^{(0)})+\operatorname{TV}(\mathbf{x}^{(0)},\mathbf{x}^{(i)})\leq 2\sqrt{T}\exp(-\sqrt{T}/8)+\frac{6\sqrt{2}}{\sqrt{\pi}}\cdot T^{-1/20}\leq 10\cdot T^{-1/20}\enspace,$
where the last inequality is true for $T\geq 400$, thus proving the lemma.
Using Lemma 6.2 and Lemma 6.4, we arrive at
$\Pr[\mathrm{Alg}(\mathbf{x}^{(i)},\mathbf{x}^{(0)})\in E_{i}]$ $\geq\Pr[\mathrm{err}(\mathcal{M}(\mathbf{x}^{(i)}))\leq\alpha]\cdot\Pr[\mathrm{err}(\mathcal{M}(\mathbf{x}^{(0)}))\leq\alpha]$
$\geq\big{(}\Pr[\mathrm{err}(\mathcal{M}(\mathbf{x}))\leq\alpha]-\operatorname{TV}(\mathbf{x}^{(i)},\mathbf{x})\big{)}$
×$\times\big{(}\Pr[\mathrm{err}(\mathcal{M}(\mathbf{x}))\leq\alpha]-\operatorname{TV}(\mathbf{x}^{(0)},\mathbf{x})\big{)}$
$\geq(3/4-\Delta^{\prime})(3/4-\Delta)\geq 1/2\enspace.$
where the last inequality holds for $T$ greater than an absolute constant. Next observe that $\mathbf{x}^{(i)}$ and $\mathbf{x}^{(0)}$ are $k$-neighboring with probability 1, and so we have $\Pr[\mathrm{Alg}(\mathbf{x}^{(i)},\mathbf{x}^{(0)})\in E_{i}]\leq e^{k\varepsilon}\Pr[\mathrm{Alg}(\mathbf{x}^{(0)},\mathbf{x}^{(0)})\in E_{i}]$, and therefore $\Pr[\mathrm{Alg}(\mathbf{x}^{(0)},\mathbf{x}^{(0)})\in E_{i}]\geq e^{-k\varepsilon}/2$. Since all $E_{j}$ are disjoint, we get
$1\geq\sum_{j=1}^{m}\Pr[\mathrm{Alg}(\mathbf{x}^{(0)},\mathbf{x}^{(0)})\in E_{j}]\geq m\cdot e^{-k\varepsilon}/2\enspace,$
and therefore
$k\geq\varepsilon^{-1}\log(m/2)\enspace.$
Recalling that $m=T/B=\sqrt{T}$, we thus have $k=\Omega(\varepsilon^{-1}\log T)$ and therefore $\alpha=k/4=\Omega(\varepsilon^{-1}\log T)$.
@@ 7 Discussion
While there are no known fundamental reasons ruling out asymptotic improvements in the mean squared error for $\varepsilon$-DP – it might be possible to match the $\Omega(\log(T)^{2})$ lower bound of [25] – there have been no asymptotic improvements since the binary tree mechanism [16, 5]. This state of affairs motivates exploring how much leading constants can be improved using the factorization mechanism framework. Algorithm 2 falls into the framework, is easy to implement, efficient, and offers the best known error-scaling to date.
Our paper also reveals that using Laplace noise improves the best possible $(\varepsilon,\delta)$-differentially private factorization mechanisms based on Gaussian noise whenever $\delta$ is sufficiently small. An interesting future direction is to investigate, for fixed $\varepsilon,\delta$ and a given factorization: which noise distribution achieves optimal error? It seems unlikely that Laplace or Gaussian noise is always best – indeed, Vinterbo [38] has results of this flavor in more restricted settings.
Finally, while we show a lower bound for random inputs that matches the best lower bound known for worst-case inputs, it is still possible that random inputs are easier and allow lower error than what can be achieved for worst-case inputs.
Acknowledgement. We are grateful to Monika Henzinger and Jalaj Upadhyay for discussions on continual counting that inspired part of this work.